# Direct $N$-body simulations of rotating and extremely massive Population III star clusters

### 31 Mar 2026 v1

@@ 5 Summary and conclusions
All simulations produce (IM)BHs below, inside, and above the PISNe mass gap (Figs. 7 and 1). The most massive IMBH has 943.19 $\mathrm{M}_{\odot}$ (from NoK1.2; Fig. 13). IMBHs above the gap can form via single-star core collapse; elsewhere they require binary channels — primarily MS-BH hyperbolic collisions, CHeBs/ShHeBs coalescence with BHs, or MS-MS mergers in massive primordial binaries. Many BHs that cannot be produced by single star evolution instead form after a common-envelope phase and subsequent coalescence that produces a CHeB then ShHeB star, which then core-collapses into a BH.
@@
With one realization per setup, we separate our findings into (i) dynamically robust trends dominated by continuous quantities or large-number counts, including global structural evolution and time-scales (core-collapse time and post-collapse self-similar expansion; Figs. 3 and 8), angular-momentum transport signatures (Fig. 11), and total mass loss / total escaper numbers (Table 5); (ii) potentially stochastic trends dominated by small-number event statistics: specific collision/coalescence subtypes and compact-object channels (e.g. BHBH collision counts; Table 4), rare escaper subtypes (e.g. $n_{\mathrm{escCHeB,s}}$), and single exceptional events (e.g. the lone BHNS merger).
@@
Cluster evolution therefore reflects a competition: gravothermal-gravogyro contraction boosts early merger rates, while the loss of massive objects due to excess angular momentum suppresses subsequent IMBH growth and BH mergers. Other physical effects, which may increase (IM)BH retention even for rotating clusters are: (i) The presence of a dark matter mini-halo around Pop III clusters, as considered by Wang et al. (2022) (we note that those simulations did not include rotation). However, Pop III star clusters may also form in larger dark matter halos, and would be dominated by their baryons after cooling; see discussion in Kimm et al. (2016). (ii) Gas infall could add more mass to the central regions, driving stronger contraction than predicted by gravothermal-gravogyro evolution. Such a contraction may trigger collapse of the central rotating IMBH subsystem and produce very massive seed BHs (see Kroupa, 2020).
@@ Appendix A Additional figures
Figure 6: Five panels (top to bottom): the number of total binaries $n_{\mathrm{b}}$, compact binaries $n_{\mathrm{cb}}$, BHBH $n_{\mathrm{BHBH}}$, BH-star $n_{\mathrm{BHstar}}$, and NS binaries $n_{\mathrm{NSb}}$ for eight simulations.
Figure 7: Plots showing the initial-final-mass relation (IFMR): the BH mass $m_{\mathrm{BH}}$ versus the progenitor ZAMS star mass $m_{\mathrm{ZAMS}}$ for all eight simulations. Columns from left to right show increasing rotation $\omega_{0}=0.0,\,0.6,\,1.2,\,1.8$. K runs use a white background; NoK runs are shaded light grey. Blue lines separate the progenitor-mass regimes, as annotated in red. Orange lines enclose the pair-instability mass gap ($\sim$64-161 $\mathrm{M}_{\odot}$, Woosley & Heger 2021). IMBHs appear below, within and above the PISN gap.
Figure 8: Plot showing the Lagrange radii $r_{\mathrm{Lagr}}$ and the average mass $M_{\mathrm{av}}$ within spheres that contain 1%, 5%, 10%, 30%, 50%, and 90% of the total cluster mass at the current simulation time-step for up to 500 Myr. Time is shown on a logarithmic scale to highlight the cluster’s rapid early evolution. Each column represents one rotational parameter $\omega_{0}$ of the rotating King model ($\omega_{0}$ = 0.0, 0.6, 1.2, 1.8). Models with K are plotted on a white background; models without GR kicks (NoK) are shaded light grey.
@@
Figure 10: Plots showing cumulative counts of escaping binary members $n_{\mathrm{esc,binarymembers}}$ and of binary components: $n_{\mathrm{escMS,b}}$, $n_{\mathrm{escCHeB,b}}$, $n_{\mathrm{escShHeB,b}}$, $n_{\mathrm{escBH,b}}$. K: solid curves; NoK: dash-dotted curves.
@@
Figure 13: Primary vs. secondary mass ($m_{1}>m_{2}$) for collisions from all eight simulations: MS–MS (orange), BH–MS (purple), BH–BH (black). Columns show $\omega_{0}=0.0,\,0.6,\,1.2,\,1.8$ left-to-right. K models are shaded white; NoK models are shaded grey. The red line denotes equal mass ratio ($q=1$). Counts in Table 4.
Figure 14: Plot of GRK velocity $v_{\mathrm{GRk}}$ versus mass ratio $q$ of the participating compact objects. All but one event are BHBH mergers; the exception is a BHNS merger, marked by an extreme $q$ and low $v_{\mathrm{GRk}}$. The shaded part of the plot corresponds to the interval 26.9 – 159.4 $\mathrm{kms}^{-1}$, the escape speeds range in cluster cores among all models.
@@ Appendix B Tables of object counts
All tables in this section report raw integer counts from a single realization per setup. For any event or object-count $N$, the intrinsic counting uncertainty is approximately Poisson, with a reference 68% ($\approx 1\sigma$) error $\sigma_{N}\simeq\sqrt{N}$ (relative error $\simeq 1/\sqrt{N}$). For small $N$, confidence intervals are non-Gaussian and can be asymmetric; therefore, differences of a few events between models should be interpreted cautiously when NN is in the single-digit or few-tens regime.
@@
Table 5: Escaping-object counts at 500 Myr in the eight simulations. [^13] The upper block gives escaping single stars: total $n_{\mathrm{esc,singles}}$, MS $n_{\mathrm{escMS,s}}$, CHeB $n_{\mathrm{escCHeB,s}}$, ShHeB $n_{\mathrm{escShHeB,s}}$, NS $n_{\mathrm{escNS,s}}$, and BH $n_{\mathrm{escBH,s}}$; the corresponding escaping single-star mass is listed as $M_{\mathrm{esc,singles}}$. The lower block gives escaping binary members: total $n_{\mathrm{esc,binarymembers}}$, MS $n_{\mathrm{escMS,b}}$, CHeB $n_{\mathrm{escCHeB,b}}$, ShHeB $n_{\mathrm{escShHeB,b}}$, NS $n_{\mathrm{escNS,b}}$, and BH $n_{\mathrm{escBH,b}}$. Notes on the uncertainty are given in Sect. B.
@@ C.1 Nbody6++GPU
Nbody6++GPU [^14] https://github.com/nbody6ppgpu is optimized for large-scale computing clusters by using MPI (Spurzem 1999), SIMD, OpenMP and GPU (Nitadori & Aarseth 2012; Wang et al. 2015) parallelization techniques. In combination with the Kustaanheimo-Stiefel (KS) regularization (Stiefel & Kustaanheimo 1965) for binary stars and close encounters, the Hermite scheme with hierarchical block time-steps (McMillan 1986; Hut et al. 1995; Makino 1991, 1999) and the Ahmad-Cohen (AC) neighbour scheme (Ahmad & Cohen 1973), two-body and chain regularization (Mikkola & Tanikawa 1999b, a; Mikkola & Aarseth 1998) the code allows for star cluster simulations of realistic size with high accuracy, including binaries.
The AC scheme divide the gravitational forces acting on it into the regular component, originating from distant stars, and an irregular part, originating from nearby stars (”neighbours”). Regular forces, efficiently accelerated on the GPU, are updated in larger regular time-steps, while neighbour forces are much more fluctuating and need updates in shorter time intervals. Since neighbour numbers are usually small compared to the total particle number, their implementation on the CPU using OpenMP (Wang et al. 2015) provides the best overall performance. Post-Newtonian dynamics of relativistic binaries is currently still using the orbit-averaged Peters & Matthews formalism (Peters & Mathews 1963; Peters 1964), as described e.g. in Di Carlo et al. (2019, 2020a, 2020b, 2021); Rizzuto et al. (2021, 2022); Arca-Sedda et al. (2021), but many upgrades were made in Arca Sedda et al. (2024) and this work.
@@
C.2 McLuster and Fopax: Initialisation of rotating star clusters
@@
    Then, we generate a 2-D Fokker-Planck initial models using Fopax (Einsel & Spurzem 1999; Kim et al. 2002, 2004, 2008). The code produces a 2-D mesh based output of density $\rho$ and velocity dispersions $\sigma$ as a function of $r$ and $z$ (as cylindrical coordinates) based on the rotating King model $f(E,J_{\mathrm{z}})$ that are characterised by a pair of parameters $(W_{0},\omega_{0})$: $f_{\mathrm{rk}}\penalty 10000\ \propto\penalty 10000\ \left(\mathrm{e}^{\beta E}-1\right)\times\penalty 10000\ \mathrm{e}^{-\beta\Omega_{0}J_{\mathrm{z}}},$ where $\beta=1/(m\sigma_{\mathrm{c}}^{2})$ and the dimensionless rotation parameter is given by $\omega_{0}=\sqrt{9/4\times\pi Gn_{\mathrm{c}}}\times\Omega_{0}$. $\sigma_{\mathrm{c}}$, $n_{\mathrm{c}}$ are central one-dimensional velocity dispersion and the central density, respectively. Potential-density pairs (e.g. Binney & Tremaine 2008) for these models are created by relating $\beta$ to the King parameter $W_{0}$ via $W_{0}=\beta m(\psi-\psi_{\mathrm{t}})$, where $\psi$ and $\psi_{\mathrm{t}}$ are the central King potential and the King potential at the truncation radius, which is normally characterized by $r_{\mathrm{tid}}$, the tidal radius (for numerical and computational methods of this step, see also Henyey et al. 1959; Cohn 1979; Spurzem 1994, 1996).
@@ C.3 SSE, BSE and Pop III stellar evolution
The role of stellar mass is extremely important, especially at larger masses. For stars less massive than 8 $\mathrm{M}_{\odot}$, we use the classical stellar evolution recipes of Hurley et al. (2000, 2002) with recent upgrades in Kamlah et al. (2022a). Adjustments for these recipes are applied where extremely metal-poor (EMP; Pop III) stars evolve differently due to the absence of heavy elements. Therefore, we assume EMP stars have negligible mass loss from winds and pulsations (Nakauchi et al. 2020) at extremely low metallicities. Since we use top-heavy initial mass functions, thus using numerous stars with $M/\mathrm{M}_{\odot}>8$, the classical stellar evolution recipes fail to model Pop III stars. In this paper, we present the implementation into Nbody6++GPU of the extended fitting formulae derived from the fitting to 1-D HOSHI stellar evolution models (Takahashi et al. 2016, 2018, 2019; Yoshida et al. 2019) of extremely metal-poor stars (Tanikawa et al. 2020, 2021b, 2021a; Hijikawa et al. 2021). At the metallicity that we adopt in this work, $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, the fitting formulae by Tanikawa et al. (2020) are generally valid from $8$ to $1280$ $\mathrm{M}_{\odot}$. We adopt the extrapolation procedure of Wang et al. (2022) for stars outside this range. Extrapolation to stellar mass beyond $1280$ $\mathrm{M}_{\odot}$ till several $10^{3}$ $\mathrm{M}_{\odot}$ can be done with these formulae, because no abrupt changes in the fitting are expected for these large masses (e.g., stellar luminosity is almost proportional to stellar mass and the star lifetime is quite similar at these masses) Extrapolation is more difficult below $8\penalty 10000\ \mathrm{M}_{\odot}{}$, where stellar mass has significantly greater effects on stellar evolution. For stars with masses below $8\penalty 10000\ \mathrm{M}_{\odot}{}$, we adopt the evolutionary prescriptions of Hurley et al. (2000) at the lowest available metallicity of $Z/\mathrm{Z}_{\odot}{}=10^{-2}$ Since only a small fraction of stars fall into this mass range under our initial conditions (see Table 1), this choice does not significantly affect the accuracy of our models. To summarize, besides Wang et al. (2022), we are among the first groups to combine direct N-body simulation with full Pop III stellar evolution from Tanikawa et al. (2020).
@@
We briefly highlight the implementations in the evolution of the stars for $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, imported from Tanikawa et al. (2020). Normally, the stellar evolution is divided into five distinct phases in chronological order: Main Sequence (MS), Hertzsprung gap (HG), Core-Helium Burning (CHeB), Shell-Helium Burning (ShHeB) and the remnant phases of either NSs or BHs. In addition, Tanikawa et al. (2020) define the blue supergiant (BSG) and red supergiant (RSG) phases that relate to the stellar surface, i.e. the effective temperature $T_{\mathrm{eff}}$, of the stars. Stars with $\mathrm{log}_{10}(T_{\mathrm{eff}}/K)\geq 3.65$ are classified as BSGs. Likewise, stars with $\mathrm{log}_{10}(T_{\mathrm{eff}}/K)<3.65$ are classified with RSGs.
@@
We refer readers to Tanikawa et al. (2020) for detailed stellar evolution parameters, or check our code publicly available on GitHub [^15] https://github.com/nbody6ppgpu/Nbody6PPGPU-beijing/tree/EMP. Using a time step parameter $\eta=0.01$ (Spurzem & Kamlah 2023), the local error scales as $\sqrt{\eta}$. The relative energy error is $\sim 10^{-8}$ per N-body time unit ($\sim 10^{4}$ yr). Due to the uncorrelated nature of these errors, the global accumulation is slow, reaching only $\sim 10^{-6}$ after 100 Myr. Given the system’s chaotic nature, this precision is sufficient. Moreover, iterative Hermite schemes can further improve energy conservation by achieving near time-symmetry (e.g., Kokubo et al. 1998; Glaschke et al. 2014; Amaro-Seoane et al. 2014; Spurzem & Kamlah 2023).
C.4 General relativistic merger recoil kicks (GRK)
The latest studies of IMBH growth (e.g. Di Carlo et al. 2019, 2020a, 2020b, 2021; Rizzuto et al. 2021, 2022) and star cluster dynamics with Nbody6++GPU (e.g. Kamlah et al. 2022a, b) do not include a general relativistic (GR) merger recoil treatment (in addition to missing PN terms). Arca-Sedda et al. (2021) included GRK via a posteriori analysis, but this does not replace a fully self-consistent treatment during the simulation. The updated Nbody6++GPU presented in this paper and Nbody7 (Aarseth 2012; Banerjee et al. 2020; Banerjee 2021) contain a proper treatment of such kicks (also see Arca Sedda et al. 2024). These velocity kicks depend on spins and mass ratio, due to asymmetric GW radiation during the final inspiral and merger. Numerical relativity (NR) models (Campanelli et al. 2007; Rezzolla et al. 2008; Hughes 2009; van Meter et al. 2010; Jiménez-Forteza et al. 2017) have been used to formulate semi-analytic descriptions for MOCCA and Nbody code family (Morawski et al. 2018, 2019; Banerjee 2021; Belczynski & Banerjee 2020; Arca-Sedda et al. 2021; Banerjee 2022). The smaller the spin (e.g., Fuller model, (nearly) non-spinning BHs), the larger the kick velocity. The larger the mass ratio, the smaller the kick velocity (Morawski et al. 2018, 2019). In the other extreme case with non-aligned natal spins and small mass ratios, the asymmetry in the GW may produce GR merger recoils that reach of order $1000\penalty 10000\ \mathrm{km\penalty 10000\ s}^{-1}{}$ (Baker et al. 2008; van Meter et al. 2010).

### 13 Apr 2026 v2

@@ 5 Summary and conclusions
All simulations produce (IM)BHs below, inside, and above the PISN mass gap (Figs. 7 and 1). The most massive IMBH has 943.19 $\mathrm{M}_{\odot}$ (from NoK1.2; Fig. 13). Single-star core collapse can form IMBHs above the gap; elsewhere IMBH formation requires binary channels — primarily MS-BH hyperbolic collisions, CHeB or ShHeB coalescence with BHs, or MS-MS mergers in massive primordial binaries. Many BHs that cannot be produced by single star evolution instead form after a common-envelope phase and subsequent coalescence that produces a CHeB then ShHeB star, which then core-collapses into a BH.
@@
With one realization per set-up, we separated our findings into (i) dynamically robust trends dominated by continuous quantities or large-number counts, including global structural evolution and timescales (core-collapse time and post-collapse self-similar expansion; Figs. 3 and 8), angular-momentum transport signatures (Fig. 11), and total mass loss as well as total escaper numbers (Table 5); (ii) potentially stochastic trends dominated by small-number event statistics, such as specific collision or coalescence subtypes and compact-object channels (e.g. BHBH collision counts; Table 4), rare escaper subtypes (e.g. $n_{\mathrm{escCHeB,s}}$), and single exceptional events (e.g. the lone BHNS merger).
@@
Cluster evolution therefore reflects a competition: gravothermal-gravogyro contraction boosts early merger rates, while the loss of massive objects due to excess angular momentum suppresses subsequent IMBH growth and BH mergers. Other physical effects, which may increase (IM)BH retention even for rotating clusters are: (i) The presence of a dark matter mini-halo around Pop III clusters, as considered by Wang et al. (2022, we note that those simulations did not include rotation). However, Pop III star clusters may also form in larger dark matter halos, and would be dominated by their baryons after cooling; see discussion in Kimm et al. (2016). (ii) Gas infall could add more mass to the central regions, driving stronger contraction than predicted by gravothermal-gravogyro evolution. Such a contraction may trigger collapse of the central rotating IMBH subsystem and produce very massive seed BHs (see Kroupa, 2020).
@@ Appendix A Additional figures
Figure 6: Top to bottom: Number of total binaries ($n_{\mathrm{b}}$), compact binaries ($n_{\mathrm{cb}}$), BHBHs ($n_{\mathrm{BHBH}}$), BH stars ($n_{\mathrm{BHstar}}$), and NS binaries ($n_{\mathrm{NSb}}$) for eight simulations.
Figure 7: Initial-final-mass relation (IFMR): the BH mass $m_{\mathrm{BH}}$ versus the progenitor ZAMS star mass $m_{\mathrm{ZAMS}}$ for all eight simulations. Columns from left to right show increasing rotation $\omega_{0}=0.0,\,0.6,\,1.2,\,1.8$. K runs use a white background; NoK runs are shaded light grey. Blue lines separate the progenitor-mass regimes, as annotated in red. Orange lines enclose the pair-instability mass gap ($\sim$64-161 $\mathrm{M}_{\odot}$, Woosley & Heger 2021). IMBHs appear below, within and above the PISN gap.
Figure 8: Lagrange radii ($r_{\mathrm{Lagr}}$) and average mass ($M_{\mathrm{av}}$) within spheres that contain 1%, 5%, 10%, 30%, 50%, and 90% of the total cluster mass at the current simulation time step for up to 500 Myr. Time is shown on a logarithmic scale to highlight the cluster’s rapid early evolution. Each column represents one rotational parameter $\omega_{0}$ of the rotating King model ($\omega_{0}$ = 0.0, 0.6, 1.2, 1.8). Models with K are plotted on a white background; models without GR kicks (NoK) are shaded light grey.
@@
Figure 10: Cumulative counts of escaping binary members $n_{\mathrm{esc,binarymembers}}$ and of binary components: $n_{\mathrm{escMS,b}}$, $n_{\mathrm{escCHeB,b}}$, $n_{\mathrm{escShHeB,b}}$, and $n_{\mathrm{escBH,b}}$. K: solid curves; NoK: dash-dotted curves.
@@
Figure 13: Primary vs secondary mass ($m_{1}>m_{2}$) for collisions from all eight simulations: MS–MS (orange), BH–MS (purple), BH–BH (black). Columns show $\omega_{0}=0.0,\,0.6,\,1.2,\,1.8$ left-to-right. K models are shaded white; NoK models are shaded grey. The red line denotes equal mass ratio ($q=1$). Counts in Table 4.
Figure 14: GRK velocity, $v_{\mathrm{GRk}}$, versus mass ratio, $q$, of the participating compact objects. All but one event are BHBH mergers; the exception is a BHNS merger, marked by an extreme $q$ and low $v_{\mathrm{GRk}}$. The shaded part of the plot corresponds to the interval 26.9 – 159.4 $\mathrm{kms}^{-1}$, the escape speeds range in cluster cores among all models.
@@ Appendix B Tables of object counts
All tables in this section report raw integer counts from a single realization per set-up. For any event or object-count $N$, the intrinsic counting uncertainty is approximately Poisson, with a reference 68% ($\approx 1\sigma$) error $\sigma_{N}\simeq\sqrt{N}$ (relative error $\simeq 1/\sqrt{N}$). For small $N$, confidence intervals are non-Gaussian and can be asymmetric; therefore, differences of a few events between models should be interpreted cautiously when NN is in the single-digit or few-tens regime.
@@
Table 5: Escaping-object counts at 500 Myr in the eight simulations. [^14] The upper block gives escaping single stars: total $n_{\mathrm{esc,singles}}$, MS $n_{\mathrm{escMS,s}}$, CHeB $n_{\mathrm{escCHeB,s}}$, ShHeB $n_{\mathrm{escShHeB,s}}$, NS $n_{\mathrm{escNS,s}}$, and BH $n_{\mathrm{escBH,s}}$; the corresponding escaping single-star mass is listed as $M_{\mathrm{esc,singles}}$. The lower block gives escaping binary members: total $n_{\mathrm{esc,binarymembers}}$, MS $n_{\mathrm{escMS,b}}$, CHeB $n_{\mathrm{escCHeB,b}}$, ShHeB $n_{\mathrm{escShHeB,b}}$, NS $n_{\mathrm{escNS,b}}$, and BH $n_{\mathrm{escBH,b}}$. Notes on the uncertainty are given in Appendix B.
@@ C.1 Nbody6++GPU
Nbody6++GPU [^15] https://github.com/nbody6ppgpu is optimized for large-scale computing clusters by using MPI (Spurzem 1999), SIMD, OpenMP and GPU (Nitadori & Aarseth 2012; Wang et al. 2015) parallelization techniques. In combination with the Kustaanheimo-Stiefel (KS) regularization (Stiefel & Kustaanheimo 1965) for binary stars and close encounters, the Hermite scheme with hierarchical block time steps (McMillan 1986; Hut et al. 1995; Makino 1991, 1999) and the Ahmad-Cohen (AC) neighbour scheme (Ahmad & Cohen 1973), two-body and chain regularization (Mikkola & Tanikawa 1999b, a; Mikkola & Aarseth 1998) the code allows for highly accurate simulations to be made of star clusters of a realistic size, including binaries.
The AC scheme divide the gravitational forces acting on it into the regular component, originating from distant stars, and an irregular part, originating from nearby stars (”neighbours”). Regular forces, efficiently accelerated on the GPU, are updated in larger regular time steps, while neighbour forces are much more fluctuating and need updates in shorter time intervals. Since neighbour numbers are usually small compared to the total particle number, their implementation on the CPU using OpenMP (Wang et al. 2015) provides the best overall performance. Post-Newtonian dynamics of relativistic binaries is currently still using the orbit-averaged Peters & Matthews formalism (Peters & Mathews 1963; Peters 1964), as described e.g. in Di Carlo et al. (2019, 2020a, 2020b, 2021); Rizzuto et al. (2021, 2022); Arca-Sedda et al. (2021), but many upgrades were made in Arca Sedda et al. (2024) and this work.
@@
C.2 McLuster and Fopax: Initialization of rotating star clusters
@@
    Then, we generate a 2-D Fokker-Planck initial models using Fopax (Einsel & Spurzem 1999; Kim et al. 2002, 2004, 2008). The code produces a 2-D mesh based output of density $\rho$ and velocity dispersions $\sigma$ as a function of $r$ and $z$ (as cylindrical coordinates) based on the rotating King model $f(E,J_{\mathrm{z}})$ that are characterized by a pair of parameters $(W_{0},\omega_{0})$: $f_{\mathrm{rk}}\penalty 10000\ \propto\penalty 10000\ \left(\mathrm{e}^{\beta E}-1\right)\times\penalty 10000\ \mathrm{e}^{-\beta\Omega_{0}J_{\mathrm{z}}},$ where $\beta=1/(m\sigma_{\mathrm{c}}^{2})$ and the dimensionless rotation parameter is given by $\omega_{0}=\sqrt{9/4\times\pi Gn_{\mathrm{c}}}\times\Omega_{0}$. $\sigma_{\mathrm{c}}$, $n_{\mathrm{c}}$ are central one-dimensional velocity dispersion and the central density, respectively. Potential-density pairs (e.g. Binney & Tremaine 2008) for these models are created by relating $\beta$ to the King parameter $W_{0}$ via $W_{0}=\beta m(\psi-\psi_{\mathrm{t}})$, where $\psi$ and $\psi_{\mathrm{t}}$ are the central King potential and the King potential at the truncation radius, which is normally characterized by $r_{\mathrm{tid}}$, the tidal radius (for numerical and computational methods of this step, see also Henyey et al. 1959; Cohn 1979; Spurzem 1994, 1996).
@@ C.3 SSE, BSE, and Pop III stellar evolution
The role of stellar mass is extremely important, especially at larger masses. For stars less massive than 8 $\mathrm{M}_{\odot}$, we use the classical stellar evolution recipes of Hurley et al. (2000, 2002) with recent upgrades in Kamlah et al. (2022a). Adjustments for these recipes are applied where extremely metal-poor (EMP; Pop III) stars evolve differently due to the absence of heavy elements. Therefore, we assume EMP stars have negligible mass loss from winds and pulsations (Nakauchi et al. 2020) at extremely low metallicities. Since we use top-heavy IMFs, thus using numerous stars with $M/\mathrm{M}_{\odot}>8$, the classical stellar evolution recipes fail to model Pop III stars. In this paper, we present the implementation into Nbody6++GPU of the extended fitting formulae derived from the fitting to 1-D HOSHI stellar evolution models (Takahashi et al. 2016, 2018, 2019; Yoshida et al. 2019) of extremely metal-poor stars (Tanikawa et al. 2020, 2021b, 2021a; Hijikawa et al. 2021). At the metallicity that we adopt in this work, $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, the fitting formulae by Tanikawa et al. (2020) are generally valid from $8$ to $1280$ $\mathrm{M}_{\odot}$. We adopt the extrapolation procedure of Wang et al. (2022) for stars outside this range. Extrapolation to stellar mass beyond $1280$ $\mathrm{M}_{\odot}$ till several $10^{3}$ $\mathrm{M}_{\odot}$ can be done with these formulae, because no abrupt changes in the fitting are expected for these large masses (e.g. stellar luminosity is almost proportional to stellar mass and the star lifetime is quite similar at these masses) Extrapolation is more difficult below $8\penalty 10000\ \mathrm{M}_{\odot}{}$, where stellar mass has significantly greater effects on stellar evolution. For stars with masses below $8\penalty 10000\ \mathrm{M}_{\odot}{}$, we adopt the evolutionary prescriptions of Hurley et al. (2000) at the lowest available metallicity of $Z/\mathrm{Z}_{\odot}{}=10^{-2}$ Since only a small fraction of stars fall into this mass range under our initial conditions (see Table 1), this choice does not significantly affect the accuracy of our models. To summarize, besides Wang et al. (2022), we are among the first groups to combine direct N-body simulation with full Pop III stellar evolution from Tanikawa et al. (2020).
@@
We briefly highlight the implementations in the evolution of the stars for $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, imported from Tanikawa et al. (2020). Normally, the stellar evolution is divided into five distinct phases in chronological order: Main Sequence (MS), Hertzsprung gap (HG), Core-Helium Burning (CHeB), Shell-Helium Burning (ShHeB) and the remnant phases of either NS or BH. In addition, Tanikawa et al. (2020) define the BSG and RSG phases that relate to the stellar surface, i.e. the effective temperature $T_{\mathrm{eff}}$, of the stars. Stars with $\mathrm{log}_{10}(T_{\mathrm{eff}}/K)\geq 3.65$ are classified as BSGs. Likewise, stars with $\mathrm{log}_{10}(T_{\mathrm{eff}}/K)<3.65$ are classified with RSGs.
@@
We refer readers to Tanikawa et al. (2020) for detailed stellar evolution parameters, or check our code publicly available on GitHub [^16] https://github.com/nbody6ppgpu/Nbody6PPGPU-beijing/tree/EMP. Using a time step parameter $\eta=0.01$ (Spurzem & Kamlah 2023), the local error scales as $\sqrt{\eta}$. The relative energy error is $\sim 10^{-8}$ per N-body time unit ($\sim 10^{4}$ yr). Due to the uncorrelated nature of these errors, the global accumulation is slow, reaching only $\sim 10^{-6}$ after 100 Myr. Given the system’s chaotic nature, this precision is sufficient. Moreover, iterative Hermite schemes can further improve energy conservation by achieving near time-symmetry (e.g. Kokubo et al. 1998; Glaschke et al. 2014; Amaro-Seoane et al. 2014; Spurzem & Kamlah 2023).
C.4 General relativistic merger recoil kicks (GRKs)
The latest studies of IMBH growth (e.g. Di Carlo et al. 2019, 2020a, 2020b, 2021; Rizzuto et al. 2021, 2022) and star cluster dynamics with Nbody6++GPU (e.g. Kamlah et al. 2022a, b) do not include a GR merger recoil treatment (in addition to missing PN terms). Arca-Sedda et al. (2021) included GRKs via a posteriori analysis, but this does not replace a fully self-consistent treatment during the simulation. The updated Nbody6++GPU presented in this paper and Nbody7 (Aarseth 2012; Banerjee et al. 2020; Banerjee 2021) contain a proper treatment of such kicks (also see Arca Sedda et al. 2024). These velocity kicks depend on spins and mass ratio, due to asymmetric GW radiation during the final inspiral and merger. Numerical relativity (NR) models (Campanelli et al. 2007; Rezzolla et al. 2008; Hughes 2009; van Meter et al. 2010; Jiménez-Forteza et al. 2017) have been used to formulate semi-analytic descriptions for MOCCA and Nbody code family (Morawski et al. 2018, 2019; Banerjee 2021; Belczynski & Banerjee 2020; Arca-Sedda et al. 2021; Banerjee 2022). The smaller the spin (e.g. Fuller model, (nearly) non-spinning BHs), the larger the kick velocity. The larger the mass ratio, the smaller the kick velocity (Morawski et al. 2018, 2019). In the other extreme case with non-aligned natal spins and small mass ratios, the asymmetry in the GW may produce GR merger recoils that reach of order $1000\penalty 10000\ \mathrm{km\penalty 10000\ s}^{-1}{}$ (Baker et al. 2008; van Meter et al. 2010).